In a Box I have (2 red balls, 3 white balls, and 5 green balls). What is the probability of picking the same color twice? Please show me how to calculate this.

Answer by stanbon(48569)

In a Box I have (2 red balls, 3 white balls, and 5 green balls). What is the probability of picking the same color twice? Please show me how to calculate this.
------------------
P(2red or 2white or 2green) = P(2red) + P(2white) + P(2green)
= 2C2/10C2 + 3C2/10C2 + 5C2/10C2
= (1 + 3 + 10)/45
= 14/45
================
Cheers,
Stan H.